OpenStudy (anonymous):

simplyfing radical expressions: how do you find x^3-x^2-42x ------------ 2x^2-20x+42

hero (hero):

:/

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

wait it's not radical it's rational

OpenStudy (anonymous):

rational expressions

OpenStudy (anonymous):

i learned when it's x^2 but not x^3

OpenStudy (anonymous):

i mastered radical expressions by the way

OpenStudy (akshay_budhkar):

take the x common in the numerator you get a quadratic equation which is what you know

OpenStudy (anonymous):

by everything?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

\[\large{\frac{x^3-x^2-42x}{2x^2-20x+42}= \frac{x(x^2-x-42)}{2x^2-20x+42}}\]

OpenStudy (anonymous):

oh i see, thanks!

hero (hero):

Of course it can be simplified further

OpenStudy (anonymous):

yes^

hero (hero):

neverforgetvivistee should post her solution on here to make sure it is right

OpenStudy (anonymous):

yes^

OpenStudy (anonymous):

i got x(x+6)/2(x+3)

OpenStudy (anonymous):

wait no it's x-3

OpenStudy (anonymous):

x(x+6)/2(x-3)

OpenStudy (anonymous):

yeah thats right

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

i'll master rationals soon

OpenStudy (anonymous):

GoodLuck ;)

OpenStudy (anonymous):

thanks everyone